Find z0 such that P(−z0 < z < z0) = 0.6. (Round your answer to two...

Find

z0

such that

P(−z0 < z < z0) = 0.6.

(Round your answer to two decimal places.)

z0 =

What percentile does

−z0

represent? (Round your answer to the nearest whole number.)

th percentile

What percentile does

z0

represent? (Round your answer to the nearest whole number.)

th percentile
Answer #1

Solution :
P(-z0 < Z < z0) = 0.6

P(Z < z0) - P(Z < z0) = 0.6
2P(Z < z0) - 1 = 0.6
2P(Z < z0) = 1 + 0.6
2P(Z < z0) = 1.6
P(Z < z0) = 1.6 / 2 = 0.8
P(Z < 0.84) = 0.8
z0 = 0.84

−z0 represent minus 84th percentile

z0 represent 84th percentile .